my question states :
Let 5x - 6y = 30
a)graph
b)what are x and y intercepts
c)what is slope
d)put this in slope-intercept form

Answer by rfer(16322)

B)
x int is when y=0
y int is when x=0
------------------
5x-6y=30
5x-6(0)=30
5x=30
x=6
------------
5(0)-6y=30
-6y=30
y=-5
------
C)
slope =5/6
Here is a start for you, You try the other to parts yourself.